2y' =3y+1 ; y(1) = 0 .


1
Expert's answer
2021-02-04T14:20:52-0500

2y' =3y+1 ; y(1) = 0 .


"2{dy \over dx} = 3y +1"


"\int{2 \over (3y +1)}dy=\int dx"


"2\int{1 \over (3y +1)}dy=\int dx"


"{2 \over 3} ln(3y +1)= x + c"


"ln(3y +1)={3 \over 2}x+c"


"3y +1= e^{{3 \over 2}x + c}"


"3y + 1= e^{{3 \over 2}x}. e^c"


"e^c= A"


"3y + 1=A e^{{3 \over 2}x}"


where A and c are constants


the initial conditionw given are y(0)=1


"3(0) +`1=Ae^{{3 \over 2}}"


"A={1 \over e^{{3 \over 2}}}= 0.22"